Exercise 4-26 (Algorithmic) (LO. 4)

Determine the taxable amount of Social Security benefits for the following situations.

If required, round your answers to the nearest dollar. If an amount is zero, enter "0".

a. Erwin and Eleanor are married and file a joint tax return. They have adjusted gross income of $42,400, no tax-exempt interest, and $14,840 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.

b. Assume Erwin and Eleanor have adjusted gross income of $17,400, no tax-exempt interest, and $19,140 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.

c. Assume Erwin and Eleanor have adjusted gross income of $110,500, no tax-exempt interest, and $16,575 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.
Answer #1

a. The Erwin and Eleanor must include $7,420 of the Social Security benefit in gross income. This works out as the lesser of the following:

  1. 0.50 ($14,840 Social Security benefits) = $7,420
  2. 0.50 [$42,400 + 0.50($14,840) - 32,000] = 0.50 ($17,820) = $8,910

b. The Erwin and Eleanor must include $0 of the benefits in gross income. This works out as the lesser of the following:

  1. 0.50 ($19,140 Social Security benefits) = $9,570
  2. 0.50 [$17,400 + 0.50($19,140) - $32,000] = 0.50 * -$5,030 = $0 .None of the benefits would be taxable as the result is not a positive number.

c. The Erwin and Eleanor include Social Security benefits will be $14,088.75. Include lesser of the following:

  1. 0.85 ($16,575 Social Security benefits) = $14,088.75
  2. Sum of:

i. 0.85 [$110,500 + 0.50($16,575) - $44,000 second base amount] = 0.85 (74,787.5) = $63,569.38

ii. Lesser of:

  • 0.50 ($16,575 Social Security benefits) = $8,287.5
  • 0.50 [$110,500 + 0.50($16,575) - $32,000] = 0.50 ($86,787.5) = $43,393.75

The sum equals to $63,569.38 + $8,287.5 = $71,856.88

Therefore, 85% of the Social Security benefits $14,088.75 is included in the couple's gross income.

Please up vote. Thanks!